# From “The Cliff” to “Virgil”: Mapping the Spectral Diversity of Little Red Dots with JWST/NIRSpec

### v1

@@ 6 Discussion: Toward a Unified View of the LRD Population
In the two-component model, the evolution toward bluer [0.30.9 m] colors is also accompanied by an increase in dust attenuation of the BH⋆ component. This helps explain the weaker optical curvature and higher incidence of power-law–like continua among bluer LRDs. Although greater BH⋆ attenuation in bluer galaxies might seem counterintuitive, it is reasonable if the host galaxies are more evolved, with higher metallicities and dust content, favoring stronger local attenuation near the BH⋆. This prediction also aligns with MIRI-based findings that bluer LRDs tend to exhibit redder NIR colors (e.g., rest-frame 1–3 m; pgp23b; barro24), potentially indicative of warm dust emission near the BH. In contrast, the reddest LRDs show flatter NIR SEDs, lacking strong dust re-emission features (e.g., degraaff25; setton24), reinforcing the idea that they are comparatively dust-free.
@@
7 Summary
This paper presents a detailed analysis of a large, uniformly selected sample of Little Red Dots (LRDs) across six JWST deep fields. Using an exhaustive color selection, we identify nearly 1000 photometric LRDs, of which 157 have publicly available NIRSpec/PRISM spectra in the Dawn JWST Archive (DJA) compilation. Within this spectroscopic sample, 118 LRDs have high continuum S/N in the rest-UV, enabling robust measurements of UV slopes and rest-frame UV-to-optical colors. We use this dataset to quantify the prevalence of broad-line AGN features, characterize the continuum SED shapes among LRDs, and explore correlations between color, SED type, emission line widths, line ratios, and the presence of absorption features.
@@ 7 Summary
Among photometric LRDs with NIRSpec grating spectra at , 50–70% exhibit broad emission lines. These represent lower limits, as nearly all show clear emission lines, and the identification of broad components is limited by depth and sensitivity.
The overall LRD population spans nearly 4 magnitudes in [0.30.9 m] color, indicating broad diversity in intrinsic properties. Most NIRSpec LRDs have blue UV slopes with a median and red UV-to-optical colors with median [0.30.9 m] . However, the reddest 18% ([0.30.9 m] ) have significantly redder slopes, reaching .
@@
Emission-line ratios and equivalent widths also correlate with [0.30.9 m] : the H +[N ii]/H ratio increases with color (reaching 20 in the reddest), while [O iii]5007/H decreases. BH⋆ candidates show the lowest values (1), and bluer LRDs reach up to 8. The equivalent widths of [O iii]5007and H +[O iii]50074364 decline toward redder colors, suggesting narrow-line emission becomes more prominent in bluer LRDs, consistent with increased host galaxy contribution.
@@
The wavelength where the SED transitions from flat UV to steep optical shifts with color: in redder LRDs ([0.30.9 m] ), the inflection occurs near the Balmer limit; in bluer LRDs, it shifts progressively redward, beyond [O iii] and toward H.
@@ Acknowledgments
We thank G. Brammer and the Dawn JWST Archive (DJA) for making high-level data products publicly available. DJA is an initiative of the Cosmic Dawn Center, which is funded by the Danish National Research Foundation un- der grant No. 140. PGP-G acknowledges support from grant PID2022-139567NB-I00 funded by Spanish Ministerio de Ciencia, Innovación y Universidades MCIU/AEI/10.13039/501100011033, FEDER Una manera de hacer Europa. JSD acknowledges the support of the Royal Society via the award of a Royal Society Research Professorship. This work has made use of the Rainbow Cosmological Surveys Database, which is operated by the Centro de Astrobiología (CAB), CSIC-INTA, partnered with the University of California Observatories at Santa Cruz (UCO/Lick, UCSC).

### v2

@@ VI.2 Is there any Dust in LRDs?
Although a higher BH⋆ attenuation in bluer galaxies may seem counterintuitive, it is plausible if these systems host more evolved galaxies with higher metallicities and dust content, leading to stronger local attenuation near the BH⋆. Consistent with this picture, MIRI observations indicate that bluer LRDs exhibit redder rest-frame near-infrared colors (e.g., 1–3 m; pg24a; barro25), suggestive of warm dust emission. In contrast, the reddest LRDs typically show flatter near-infrared SEDs with little evidence for dust re-emission (e.g., degraaff25; setton25), supporting the view that they are comparatively dust-poor.
The rest-frame [13 m] color provides a sensitive diagnostic of dust emission. Stellar populations and typical obscured AGNs without significant hot dust produce relatively mild colors ([13 m] 00.5; Figure 8 of barro25), and even BH⋆ or blackbody-like photospheric emission would yield similarly flat or negative values. By contrast, the median [13 m] color of MIRI-detected LRDs is closer to 1.5, implying the presence of warm dust emission or, alternatively, a more complex intrinsic continuum shape than captured by simple BH⋆ models.
@@
VII Summary
This paper presents a detailed analysis of a large, uniformly selected sample of Little Red Dots (LRDs) across six JWST deep fields. Using an exhaustive color selection, we identify nearly 1000 photometric LRDs, of which 191 have publicly available NIRSpec/PRISM spectra in the Dawn JWST Archive (DJA) compilation. Within this spectroscopic sample, 118 LRDs have high continuum S/N in the rest-UV, enabling robust measurements of UV slopes and rest-frame UV-to-optical colors. We use this dataset to quantify the prevalence of broad-line AGN features, characterize the continuum SED shapes among LRDs, and explore correlations between color, SED type, emission line widths, line ratios, and the presence of absorption features.
@@ VII Summary
Among photometric LRDs with NIRSpec grating spectra at , 50% to 90% exhibit broad emission lines, depending on the field. These represent lower limits, as nearly all show clear emission lines, and the identification of broad components is limited by depth and sensitivity.
The overall LRD population spans nearly 4 magnitudes in [0.30.9 m] rest-frame color, indicating broad diversity in intrinsic properties. Most NIRSpec LRDs have blue UV slopes with a median and red UV-to-optical colors with median [0.30.9 m] . However, the reddest 18% ([0.30.9 m] ) have significantly redder slopes, reaching .
@@
Emission-line ratios and equivalent widths also correlate with [0.30.9 m]: the H +[N ii]/H ratio increases with color (reaching 20 in the reddest), while [O iii]5007/H decreases. BH⋆ candidates show the lowest values (1), and bluer LRDs reach up to 8. The equivalent widths of [O iii]5007 and H +[O iii]4364 decline toward redder colors, suggesting narrow-line emission becomes more prominent in bluer LRDs, consistent with increased host galaxy contribution.
@@
The wavelength where the SED transitions from flat UV to steep optical shits with color: in redder LRDs ([0.30.9 m] ), the inflection occurs near the Balmer limit; in bluer LRDs, it shifts progressively redward, beyond [O iii]5007 and toward H .
@@ Acknowledgments
We thank G. Brammer and the Dawn JWST Archive (DJA) for making high-level data products publicly available. DJA is an initiative of the Cosmic Dawn Center, which is funded by the Danish National Research Foundation under grant No. 140. PGP-G acknowledges support from grant PID2022-139567NB-I00 funded by Spanish Ministerio de Ciencia, Innovación y Universidades MCIU/AEI/10.13039/501100011033, FEDER Una manera de hacer Europa.
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute (STScI), which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. These observations are associated with JWST programs GTO-1180, GTO-1181, GTO-1215, GTO-1286, GO-1345, GO-2198, GO-2561, DDT-2750, GO-3073, GO-4106, GO-4233, GO-5224, GO-6585, GO-6838, and DDT-6541.
JSD acknowledges the support of the Royal Society via the award of a Royal Society Research Professorship. This work has made use of the Rainbow Cosmological Surveys Database, which is operated by the Centro de Astrobiología (CAB), CSIC-INTA, partnered with the University of California Observatories at Santa Cruz (UCO/Lick, UCSC). This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101148925.